OpenStudy (anonymous):

The question explains that you would like to give away candy only half of the time. You have also wanted to ive away two different kinds of candy to the people who win equal amounts, one small and one large. I need the probability of each contestant losing, winning a large peice of candy, and winning a small piece of candy.

OpenStudy (apoorvk):

can you type out the exact question please? this seems very ambiguous.

OpenStudy (anonymous):

If you win half the time and then distribute two options equally you can determine the odds. Your odds of wining are half the time or 1/2. If you win your odds of getting one candy or the other are again half the time or 1/2. In order to produce the probablity of reaching each candy you need to multiply the probability of wining by the probability of each item. \[\frac{1}{2} \times \frac{1}{2} = \frac{1}{4}\]

OpenStudy (anonymous):

Thanks!